OpenStudy (anonymous):

Find the solution to the system of equations. x+y+z=22 y+z=16 z=6 If someone could explain to me how to do this I would really appreciate it! Thank you.

OpenStudy (deoxna):

Plug in the 6 for "z" in the second equation, because then you're left with only one variable: y+6=16 so y=10 Then, plug in 6 and 10 into the first equation and solve for x: x+16=22 x=6 y=10 z=6

OpenStudy (anonymous):

first plug z=6 in the second equation and you can find y: y + 6 = 16 so y = 10 now plug z=6 and y=10 in the first equation: x + 10 + 6 = 22 x = 22-10-6 = 6 x=6,y=10,z=6

OpenStudy (anonymous):

z=6 y+6 = 16 --> y = 10 x + 10 + 6 = 22 --> x = 6